1) The lengths of pregnancies in a small rural village are normally distributed with a mean of 263 days and a standard deviation of 15 days. A distribution of values is normal with a mean of 263 and a standard deviation of 15.

What percentage of pregnancies last fewer than 295 days? P(X < 295 days) = %

2) The physical fitness of an athlete is often measured by how much oxygen the athlete takes in (which is recorded in milliliters per kilogram, ml/kg). The mean maximum oxygen uptake for elite athletes has been found to be 68 with a standard deviation of 7.8. Assume that the distribution is approximately normal.

Find the probability that an elite athlete has a maximum oxygen uptake of at most 47.7 ml/kg.
Answer #1

Solution :

Given that ,

1) mean = \mu = 263

standard deviation = \sigma = 15

P(x < 295) = P[(x - \mu ) / \sigma < (295 - 263) / 15]

= P(z < 2.13)

Using z table,

=0.9834

The percentage is = 98.34%

2) mean = \mu = 68

standard deviation = \sigma = 7.8

P(x \leq 47.7)

= P[(x - \mu ) / \sigma \leq (47.7 - 68) / 7.8 ]

= P(z \leq -2.60)

Using z table,

= 0.0047